PROBLEM 7 i got wrong Factor completely. b2 – ab – 6a2

Answer by venugopalramana(3286)

PROBLEM 7 i got wrong Factor completely. b2 – ab – 6a2
=B^2-3AB+2AB-6A^2
=B(B-3A)+2A(B-3A)
=(B-3A)(B+2A)